OpenStudy (anonymous):

Jack bought 3 slices of cheese pizza and 4 sclices of mushroom pizza for $12.50. Grace bought 3 slices of cheese pizza and 2 slices of mushroom pizza for $8.50. What is the cost of 1 slice of mushroom pizza? A) $1.50 B) $2.00 C)$3.00 D)$3.50

OpenStudy (anonymous):

If 'x' is a cheese slice, 'y' is a mushroom slice you need to solve the set of equations: (1) 3x + 4y = 12.5 (2) 3x + 2y = 8.5 And find how much is 'y' Add (-2y) to both sides of equation (2) to get: 3x = 8.5 - 2y And then substitute it into the equation (1) (8.5 - 2y) + 4y = 12.5

OpenStudy (anonymous):

aliter to adamshai 's work (1) 3x + 4y = 12.5 (2) 3x + 2y = 8.5 since coefficient of x are equal just subtract eq 2 from eq 1 we get 2y=4 or y=2 the cost of 1 slice of mushroom pizza is B) $2.00